# Rhaly operators acting on Hardy and Bergman spaces

### v1

@@ Abstract.
In this article we address the question of characterizing the sequences of complex numbers $(\eta)=\{\eta_{n}\}_{n=0}^{\infty}$ whose associated Rhaly operator $\mathcal{R}_{(\eta)}$ is bounded or compact on the Hardy spaces $H^{p}$ ($1\leq p<\infty$) or on the Bergman spaces $A^{p}$ ($1\leq p<\infty$). Among other results we completely characterize those $(\eta)$ for which $\mathcal{R}_{(\eta)}$ is bounded or compact on $H^{p}$ ($1<p\leq 2$) and on $A^{p}$ ($1<p<\infty$). We also give conditions on $(\eta)$ which are either necessary or sufficient for the boundedness (compactness) of $\mathcal{R}_{(\eta)}$ on $H^{p}$ for $p=1$ and $2<p<\infty$.
In particular, we prove that if $2\leq p<\infty$ and $\eta_{n}=\operatorname{O}\left(\frac{1}{n}\right)$, then $\mathcal{R}_{(\eta)}$ is bounded on $H^{p}$. However, there exists a sequence $(\eta)$ with $\eta_{n}=\operatorname{O}\left(\frac{1}{n}\right)$ such that the operator $\mathcal{R}_{(\eta)}$ is not bounded on $H^{p}$ for $1\leq p<2$.
@@
Hardy spaces, Bergman spaces, The Cesàro operator, Rhaly operators, Mean Lipschitz spaces
@@ 1. Introduction and main results
The case when $\eta_{n}=\frac{1}{n+1}$ for all $n$ corresponds to the classical Cesàro operator $\mathcal{C}$ which is known to be bounded from $\ell^{p}$ to $\ell^{p}$ for $1<p\leq\infty$. This was proved by Hardy [21] and Landau [25] (see also [22, Theorem 326, p.239 ]). The sequences $\{\eta_{n}\}_{n=0}^{\infty}$ for which the operator $\mathcal{R}_{\{\eta_{n}\}}$ is either bounded or compact on the space $\ell^{p}$ for any $p\in(1,\infty)$ have been characterized in [19]. A characterization of the sequences $\{\eta_{n}\}$ for which $\mathcal{R}_{\{\eta_{n}\}}$ belongs to the Schatten class $\mathcal{S}^{q}(\ell^{2})$ ($1<q<\infty$) has been obtained in [4].
@@
(a) If $1<p\leq 2$, then $\mathcal{R}_{(\eta)}\in\mathcal{B}(H^{p})$  if and only if $F_{(\eta)}\in\Lambda^{p}_{1/p}$.
(b) If $2<p<\infty$ and $\mathcal{R}_{(\eta)}\in\mathcal{B}(H^{p})$, then $F_{(\eta)}\in\Lambda^{p}_{1/p}$.
@@
(a) If $1<p\leq 2$ then $\mathcal{R}_{(\eta)}\in\mathcal{K}(H^{p})$  if and only if $F_{(\eta)}\in\lambda^{p}_{1/p}$.
(b) If $2<p<\infty$ and $\mathcal{R}_{(\eta)}\in\mathcal{K}(H^{p})$, then $F_{(\eta)}\in\lambda^{p}_{1/p}$.
@@
Let $(\eta)=\{\eta_{n}\}_{n=0}^{\infty}$ be a sequence of complex numbers and $1<p<\infty$. Then:
(a) $\mathcal{R}_{(\eta)}\in\mathcal{B}(A^{p})$  if and only if $F_{(\eta)}\in\Lambda^{p}_{1/p}$.
(b) $\mathcal{R}_{(\eta)}\in\mathcal{K}(A^{p})$  if and only if $F_{(\eta)}\in\lambda^{p}_{1/p}$.
@@
The mean Lipschitz spaces $\Lambda^{p}_{1/p}$ and the “litle oh”mean Lipschitz spaces $\lambda^{p}_{1/p}$ will be properly defined in Section 2, which will be devoted to introduce a number of spaces and results which will be needed in our work. The proofs of Theorem 1 and Theorem 2 will be presented in Section 3, and Section 4 will include the proof of Theorem 3. Theorem 4 will be proved in Section 5 where we will present also a number of results regarding the operators $\mathcal{R}_{(\eta)}$ acting on $H^{1}$.
@@ 2.1. Mean Lipschitz spaces
For $1\leq p\leq\infty$ and $0<\alpha\leq 1$, we let $\Lambda^{p}_{\alpha}$ be the space of those $f\in\rm{Hol}(\mathbb{D})$ having a non-tangential limit at almost every point of $\partial\mathbb{D}$ and so that $\omega_{p}(\cdot,f)$, the integral modulus of continuity of order $p$ of the boundary values $f(e^{i\theta})$ of $f$, satisfies $\omega_{p}(\delta,f)=\operatorname{O}(\delta^{\alpha})$, as $\delta\to 0$. When $p=\infty$ we write $\Lambda_{\alpha}$ instead of $\Lambda^{\infty}_{\alpha}$. This is the usual space of order $\alpha$. More precisely, a function $f\in\rm{Hol}(\mathbb{D})$ belongs to $\Lambda_{\alpha}$ if and only if it has a continuous extension to the closed unit disc and its boundary values satisfy a Lipschitz condition of order $\alpha$.
@@
Classical results of Hardy and Littlewood (see [10] and [11, Chapter 5]) show that whenever $0<\alpha\leq 1$ and $1\leq p\leq\infty$ we have that $\Lambda^{p}_{\alpha}\subset H^{p}$ and that
@@
The corresponding “little oh”  spaces are denoted by $\lambda^{p}_{\alpha}$.
If $1<p<\infty$ and $(1/p)<\alpha\leq 1$, then $\Lambda^{p}_{\alpha}\subset\Lambda_{\alpha-{(1/p)}}$ and, hence, each function in $\Lambda^{p}_{\alpha}$ has a continuous extension to the closed unit disc [10, p. 88]. This is not true for the space $\Lambda^{p}_{1/p}$. This follows easily noticing that the function $f(z)=\log(1-z)$ is an unbounded function which belongs to $\Lambda_{1/p}^{p}$ for all $p\in(1,\infty)$. Bourdon, Shapiro, and Sledd [10] proved that $\Lambda^{p}_{1/p}\subset BMOA$ for all $p\in(1,\infty)$. This inclusion was shown to be sharp in a very strong sense in [9]. Let us remark also that the $\Lambda^{p}_{1/p}$-spaces form a nested scale of spaces
@@ 2.2. Convolution of analytic functions
$\|f\star g\|_{A^{p}_{\alpha}}\,\lesssim\|f\|_{H^{p}}\|g\|_{A^{p}_{\alpha}}.$
@@
$\|f\ast g\|_{A^{p}_{\alpha}}^{p}$ $\asymp\int_{0}^{1}(1-r)^{\alpha}M_{p}^{p}(r,f\ast g)\,dr\,\lesssim\int_{0}^{1}(1-r)^{\alpha}\|f\|_{H^{1}}^{p}M_{p}^{p}(r,g)\,dr$
$\lesssim\|f\|_{H^{p}}^{p}\int_{0}^{1}(1-r)^{\alpha}M_{p}^{p}(r,g)\,dr\,\asymp\|f\|_{H^{p}}^{p}\|g\|_{A^{p}_{\alpha}}^{p}.$
@@
2.3. “Lipschitz” Cesàro bases
Induced by the work of Pavlović [34, pp.  140-143] who constructed “smooth”  Cesàro bases, next we are going to construct “Lipschitz” Cesàro bases.
Let $\Psi:\mathbb{R}\rightarrow\mathbb{C}$ be a Lipschitz function with compact support. For simplicity, we shall assume that the support of $\Psi$ is contained in $[0,4]$. We set
$\mathcal{L}(\psi)\,=\,\sup_{s\neq t}\frac{|\Psi(t)-\Psi(s)|}{|t-s|}.$
For $n\in\mathbb{N}$ we define
$W_{n}^{\Psi}(e^{it})\,=\,\sum_{k=-\infty}^{\infty}\Psi\left(\frac{k}{n}\right)e^{ik\theta}\,=\,\sum_{k=0}^{4n}\Psi\left(\frac{k}{n}\right)e^{ik\theta}.$
Arguing as in the proof of Lemma 5.2 in p. 141 of [34] with $N=2$, we have
(2.16) $\,(1-e^{i\theta})^{2}W_{n}^{\Psi}(e^{i\theta})\,=\,\sum_{k=0}^{4n}\Psi\left(\frac{k}{n}\right)(1-e^{i\theta})^{2}e^{ik\theta}\,=\,\sum_{k=0}^{4n}\Psi\left(\frac{k}{n}\right)\sum_{m=0}^{2}\binom{2}{m}(-1)^{m}e^{i(m+k)\theta}$
$=\,$ $\sum_{m=0}^{2}(-1)^{m}\binom{2}{m}\sum_{k=0}^{4n}\Psi\left(\frac{k}{n}\right)e^{i(m+k)}\theta\,=\,\sum_{m=0}^{2}(-1)^{m}\binom{2}{m}\sum_{j=0}^{4n+2}\Psi\left(\frac{j-m}{n}\right)e^{ij\theta}$
$=\,$ $\sum_{j=0}^{4n+2}\left[\sum_{m=0}^{2}(-1)^{m}\binom{2}{m}\Psi\left(\frac{j-m}{n}\right)\right]e^{ij\theta}.$
@@
$\left|\sum_{m=0}^{2}(-1)^{m}\binom{2}{m}\Psi\left(\frac{j-m}{n}\right)\right|=$ $\left|\Psi\left(\frac{j}{n}\right)-\Psi\left(\frac{j-1}{n}\right)-\Psi\left(\frac{j-1}{n}\right)+\Psi\left(\frac{j-2}{n}\right)\right|\leq\frac{2\mathcal{L}(\psi)}{n}.$
Using this in (2.16), we obtain
(2.17) $\left|(1-e^{i\theta})^{2}W_{n}^{\Psi}(e^{i\theta})\right|\leq\frac{2\mathcal{L}(\Psi)}{n}\sum_{j=0}^{4n+2}|e^{ij\theta}|\leq\frac{2\mathcal{L}(\Psi)}{n}(4n+3)\leq 14\mathcal{L}(\Psi).$
3. Rhaly operators acting on Hardy spaces
Let us start proving that, for $1<p<\infty$, the condition $F_{(\eta)}\in\Lambda^{p}_{1/p}$ (respectively, $F_{(\eta)}\in\lambda^{p}_{1/p}$) is a necessary condition for the membership of $\mathcal{R}_{(\eta)}$ in $\mathcal{B}(H^{p})$ (respectively, in $\mathcal{K}(H^{p}))$.
Proof. Assume that $1<p<\infty$ and $\mathcal{R}_{(\eta)}\in\mathcal{B}(H^{p})$. As observed in section 1 this implies that $F_{(\eta)}$ is a well defined analytic function in $\mathbb{D}$. In fact, since $F_{(\eta)}=\mathcal{R}_{(\eta)}(1)$, $F_{(\eta)}\in H^{p}$.
For $N=2,3,\dots$, set $a_{N}=1-\frac{1}{N}$ and
(3.1) $f_{N}(z)\,=\,\frac{1}{N^{2-\frac{1}{p}}}\frac{a_{N}z}{(1-a_{N}z)^{2}}\,=\,\frac{1}{N^{2-\frac{1}{p}}}\sum_{n=1}^{\infty}na_{N}^{n}z^{n},\quad z\in\mathbb{D}.$
We have that $f_{N}\in H^{p}$ for all $N$ and
(3.2) $\|f_{N}\|_{H^{p}}\lesssim 1.$

### v2

# Rhaly operators acting on Hardy, Bergman, and Dirichlet spaces

@@ Abstract.
In this article we address the question of characterizing the sequences of complex numbers $(\eta)=\{\eta_{n}\}_{n=0}^{\infty}$ whose associated Rhaly operator $\mathcal{R}_{(\eta)}$ is bounded or compact on the Hardy spaces $H^{p}$ ($1\leq p<\infty$), on the Bergman spaces $A^{p}_{\alpha}$, and on the Dirichlet spaces $\mathcal{D}^{p}_{\alpha}$ ($1\leq p<\infty$, $\alpha>-1$). We give a number of conditions which are either necessary or sufficient for the boundedness (compactness) of $\mathcal{R}_{(\eta)}$ on these spaces. These conditions have to do with the membership in certain mean Lipschitz spaces of analytic functions of the function $F_{(\eta)}$ defined by $F_{(\eta)}(z)=\sum_{n=0}^{\infty}\eta_{n}z^{n}$ ($z\in\mathbb{D}$).
We prove that if $2\leq p<\infty$ and $\eta_{n}=\operatorname{O}\left(\frac{1}{n}\right)$, then $\mathcal{R}_{(\eta)}$ is bounded on $H^{p}$. However, there exists a sequence $(\eta)$ with $\eta_{n}=\operatorname{O}\left(\frac{1}{n}\right)$ such that the operator $\mathcal{R}_{(\eta)}$ is not bounded on $H^{p}$ for $1\leq p<2$.
We deal also with the derivative-Hardy spaces. For $p>0$ the derivative-Hardy space $S^{p}$ consists of those functions $f$, analytic in the unit disc $\mathbb{D}$, such that $f^{\prime}\in H^{p}$. We prove that if $1\leq p<\infty$ and $1<q<\infty$ then $\mathcal{R}_{(\eta)}$ is a bounded operator from $S^{p}$ into $S^{q}$ if and only if it is compact and this happens if and only if $F_{(\eta)}\in S^{q}$.
@@
Hardy spaces, Bergman spaces, Dirichlet spaces, The Cesàro operator, Rhaly operators, Mean Lipschitz spaces
@@ 1. Introduction and main results
The case when $\eta_{n}=\frac{1}{n+1}$ for all $n$ corresponds to the classical Cesàro operator $\mathcal{C}$ which is known to be bounded from $\ell^{p}$ to $\ell^{p}$ for $1<p\leq\infty$. This was proved by Hardy [23] and Landau [27] (see also [24, Theorem 326, p. 239 ]). The sequences $\{\eta_{n}\}_{n=0}^{\infty}$ for which the operator $\mathcal{R}_{\{\eta_{n}\}}$ is either bounded or compact on the space $\ell^{p}$ for any $p\in(1,\infty)$ have been characterized in [20]. A characterization of the sequences $\{\eta_{n}\}$ for which $\mathcal{R}_{\{\eta_{n}\}}$ belongs to the Schatten class $\mathcal{S}^{q}(\ell^{2})$ ($1<q<\infty$) has been obtained in [4].
@@
(a) If $1<p<\infty$ and $\mathcal{R}_{(\eta)}\in\mathcal{B}(H^{p})$ then
(1.2) $M_{p}(r,F_{(\eta)}^{\prime})\,=\,\operatorname{O}\left(\frac{\log\frac{1}{1-r}}{(1-r)^{1-\frac{1}{p}}}\right),\quad\text{ as $r\to 1$.}$
(b) If $1<p\leq 2$ and $F_{(\eta)}\in\Lambda^{p}_{1/p}$ then $\mathcal{R}_{(\eta)}\in\mathcal{B}(H^{p})$.
@@
(a) If $1<p<\infty$ and $\mathcal{R}_{(\eta)}\in\mathcal{K}(H^{p})$ then
(1.3) $M_{p}(r,F_{(\eta)}^{\prime})\,=\,\operatorname{o}\left(\frac{\log\frac{1}{1-r}}{(1-r)^{1-\frac{1}{p}}}\right),\quad\text{ as $r\to 1$.}$
(b) If $1<p\leq 2$ and $F_{(\eta)}\in\lambda^{p}_{1/p}$ then $\mathcal{R}_{(\eta)}\in\mathcal{K}(H^{p})$.
@@
Let $(\eta)\,=\,\{\eta_{n}\}_{n=0}^{\infty}$ be a sequence of complex numbers.
(iii) If $1<p<\infty$, $1<\alpha<2p-2$, and $\mathcal{R}_{(\eta)}\in\mathcal{B}(A^{p}_{\alpha})$ then (1.2) holds.
(iv) If $1<p<\infty$, $1<\alpha<2p-2$, and $\mathcal{R}_{(\eta)}\in\mathcal{K}(A^{p}_{\alpha})$ then (1.3) holds.
(v) If $\alpha>-1$, then $\mathcal{R}_{(\eta)}\in\mathcal{B}(A^{2}_{\alpha})$ if and only if $F_{(\eta)}\in\Lambda^{2}_{1/2}$.
(vi) If $\alpha>-1$, then $\mathcal{R}_{(\eta)}\in\mathcal{K}(A^{2}_{\alpha})$ if and only if $F_{(\eta)}\in\lambda^{2}_{1/2}$.
Theorem 4.
Let $(\eta)\,=\,\{\eta_{n}\}_{n=0}^{\infty}$ be a sequence of complex numbers.
(i) If $1\leq p<\infty$, $\alpha>p-2$, and $F_{(\eta)}\in\Lambda^{p}_{1/p}$ then $\mathcal{R}_{(\eta)}\in\mathcal{B}(\mathcal{D}^{p}_{\alpha})$.
(ii) If $1\leq p<\infty$, $\alpha>p-2$, and $F_{(\eta)}\in\lambda^{p}_{1/p}$ then $\mathcal{R}_{(\eta)}\in\mathcal{K}(\mathcal{D}^{p}_{\alpha})$.
(iii) If $1<p<\infty$, $1<\alpha<2p-2$, and $\mathcal{R}_{(\eta)}\in\mathcal{B}(\mathcal{D}^{p}_{\alpha})$ then (1.2) holds.
(iv) If $1<p<\infty$, $1<\alpha<2p-2$, and $\mathcal{R}_{(\eta)}\in\mathcal{K}(\mathcal{D}^{p}_{\alpha})$ then (1.3) holds.
(v) If $\alpha>0$, then $\mathcal{R}_{(\eta)}\in\mathcal{B}(\mathcal{D}^{2}_{\alpha})$ if and only if $F_{(\eta)}\in\Lambda^{2}_{1/2}$.
(vi) If $\alpha>0$, then $\mathcal{R}_{(\eta)}\in\mathcal{K}(\mathcal{D}^{2}_{\alpha})$ if and only if $F_{(\eta)}\in\lambda^{2}_{1/2}$.
@@
Finally, we are going to deal with the derivative-Hardy spaces $S^{p}$. For $p>0$, the derivative-Hardy space $S^{p}$ consists of those $f\in\rm{Hol}(\mathbb{D})$ such that $f^{\prime}\in H^{p}$. For $f\in S^{p}$, we set
$\|f\|_{S^{p}}^{p}\,=\,|f(0)|^{p}\,+\,\|f^{\prime}\|_{H^{p}}^{p}.$
In some instances the space $S^{p}$ plays the role of “the limit of $\mathcal{D}^{p}_{\alpha}$, as $\alpha\to-1^{+}$”. Corollary 1 of [8] asserts that
$\mathcal{R}_{(\eta)}\in\mathcal{B}(S^{2})\,\,\,\Leftrightarrow\,\,\,\mathcal{R}_{(\eta)}\in\mathcal{K}(S^{2})\,\,\,\Leftrightarrow\,\,\,F_{(\eta)}\in S^{2}.$
We extend this in the following theorem.
Theorem 6.
Suppose that $1\leq p<\infty$, $1<q<\infty$, and let $(\eta)=\{\eta_{n}\}_{n=0}^{\infty}$ be a sequence of complex numbers. Then the following conditions are equivalent.
(i) $\mathcal{R}_{(\eta)}\in\mathcal{B}(S^{p},S^{q})$.
(ii) $\mathcal{R}_{(\eta)}\in\mathcal{K}(S^{p},S^{q})$.
(iii) $F_{(\eta)}\in S^{q}$.
The mean Lipschitz spaces $\Lambda^{p}_{1/p}$, the “little oh”mean Lipschitz spaces $\lambda^{p}_{1/p}$, and, more generally the mean Lipschitz spaces $\Lambda(p,\omega)$ will be properly defined in Section 2, which will be devoted to introduce a number of spaces and results which will be needed in our work. The proofs of the results for $p=2$ will be presented in Section 3. Section 4 will include the proofs of the necessity statements in Theorems 1, 2, 3, and 4 for $p\neq 2$. The sufficiency statements in these theorems will be proved in Section 5. Theorem 5 will be proved in Section 6 where we will present also a number of results regarding the operators $\mathcal{R}_{(\eta)}$ acting on $H^{1}$. Finally, the proof of Theorem 6 will be presented in Section 7.
@@ 2.1. Mean Lipschitz spaces
If $f$ is a function which is analytic in $\mathbb{D}$ and has a non-tangential limit $f(e^{i\theta})$ at almost every $\xi\in\partial\mathbb{D}$, we define
$\omega_{p}(\delta,f)=\sup_{0<|t|\leq\delta}\left(\frac{1}{2\pi}\int_{-\pi}^{\pi}\left|f(e^{i(\theta+t)})-f(e^{i\theta})\right|^{p}\,d\theta\right)^{1/p},$ $\quad\delta>0,\quad\hbox{if $1\leq p<\infty$},$
$\omega_{\infty}(\delta,f)=\sup_{0<|t|\leq\delta}\left(\operatornamewithlimits{ess.sup}_{\theta\in[-\pi,\pi]}|f(e^{i(\theta+t)})-f(e^{i\theta})|\right),$ $\quad\delta>0.$
Then $\omega_{p}(.,f)$ is the integral modulus of continuity of order $p$ of the boundary values $f(e^{i\theta})$ of $f$.
Throughout the paper $\omega:[0,1]\rightarrow[0,\infty)$ will be a continuous and increasing function with $\omega(0)=0$. Then, for $1\leq p\leq\infty$, the mean Lipschitz space $\Lambda(p,\omega)$ consists of those functions $f\in H^{p}$ which satisfy
$\omega_{p}(\delta,f)=\operatorname{O}(\omega(\delta)),\quad\hbox{as $\delta\to 0$}.$
If $0<\alpha\leq 1$ and $\omega(\delta)=\delta^{\alpha}$, we shall write $\Lambda^{p}_{\alpha}$ instead of $\Lambda(p,\omega)$, that is, we set
$\Lambda^{p}_{\alpha}=\Lambda(p,\delta^{\alpha}),\quad 0<\alpha\leq 1,\quad 1\leq p\leq\infty.$
The corresponding “little oh”  spaces are denoted by $\lambda^{p}_{\alpha}$ and $\lambda(p,\omega)$.
@@
Classical results of Hardy and Littlewood (see [11] and [12, Chapter 5]) show that whenever $0<\alpha\leq 1$ and $1\leq p\leq\infty$ we have that $\Lambda^{p}_{\alpha}\subset H^{p}$ and that
@@
This result was extended in [10] where Blasco and de Souza proved that if $\omega$ satisfies the so called Dini and $b_{1}$ conditions then
$\Lambda(p,\omega)=\left\{\hbox{$f$ analytic in $\Delta$ : $M_{p}(r,f^{\prime})=\operatorname{O}\left(\frac{\omega(1-r)}{1-r}\right),$ as $r\to 1$}\right\}.$
In particular this is true for $\omega$ defined by
(2.5) $\omega(\delta)\,=\,\delta^{\alpha}\log\frac{A}{\delta},\quad 0<\delta<1,$
with $0<\alpha<1$ and $A\geq e^{1/\alpha}$. Thus, if $\omega$ is as in (2.5) with $0<\alpha<1$ and $A\geq e^{1/\alpha}$, then
(2.6) $\Lambda(p,\omega)\,=\,\left\{f\in H^{p}:M_{p}(r,f^{\prime})\,=\,\operatorname{O}\left(\frac{\log\frac{1}{1-r}}{(1-r)^{1-\alpha}}\right),\,\,\text{as $r\to 1$}\right\}.$
If $1<p<\infty$ and $(1/p)<\alpha\leq 1$, then $\Lambda^{p}_{\alpha}\subset\Lambda_{\alpha-{(1/p)}}$ and, hence, each function in $\Lambda^{p}_{\alpha}$ has a continuous extension to the closed unit disc [11, p. 88]. This is not true for the space $\Lambda^{p}_{1/p}$. This follows easily noticing that the function $f(z)=\log(1-z)$ is an unbounded function which belongs to $\Lambda_{1/p}^{p}$ for all $p\in(1,\infty)$. Bourdon, Shapiro, and Sledd [11] proved that $\Lambda^{p}_{1/p}\subset BMOA$ for all $p\in(1,\infty)$. This inclusion was shown to be sharp in a very strong sense in [9]. Let us remark also that the $\Lambda^{p}_{1/p}$-spaces form a nested scale of spaces
@@ 2.2. Convolution of analytic functions
$\|f\star g\|_{A^{p}_{\alpha}}\,\lesssim\|f\|_{H^{1}}\|g\|_{A^{p}_{\alpha}}.$
@@
$\|f\ast g\|_{A^{p}_{\alpha}}^{p}$ $\asymp\int_{0}^{1}(1-r)^{\alpha}M_{p}^{p}(r,f\ast g)\,dr\,\lesssim\int_{0}^{1}(1-r)^{\alpha}\|f\|_{H^{1}}^{p}M_{p}^{p}(r,g)\,dr\asymp\|f\|_{H^{1}}^{p}\|g\|_{A^{p}_{\alpha}}^{p}.$
@@
3. The case $p=2$
The results stated for $p=2$ are essentially contained in [20] and [8].
Indeed, Theorem 1 of [20] and Theorem A give that $\mathcal{R}_{(\eta)}\in\mathcal{B}(H^{2})$ if and only if $F_{(\eta)}\in\Lambda^{2}_{1/2}$.
Similarly, part (d) of Theorem 2 follows from Theorem 2 of [20] and the “little oh”analogue of Theorem A.
Notice that, for $f\in\rm{Hol}(\mathbb{D})$, $f(z)=\sum_{n=0}^{\infty}a_{n}z^{n}$ ($z\in\mathbb{D}$), we have that
$\|f\|_{\mathcal{D}^{2}_{\alpha}}^{2}\,\asymp\,|a_{0}|^{2}\,+\,\sum_{n=1}^{\infty}n^{1-\alpha}|a_{n}|^{2}.$
Then Theorem 3 of [8] shows that, for $\alpha>0$,
(3.1) $\mathcal{R}_{(\eta)}\in\mathcal{B}(\mathcal{D}^{2}_{\alpha})\,\,\Leftrightarrow\,\,\sum_{n=N}^{\infty}n^{1-\alpha}|\eta_{n}|^{2}\,=\,\operatorname{O}(N^{-\alpha}),\,\,\,\text{as $N\to\infty$}.$
@@ 3. The case $p=2$
$\sum_{n=1}^{N}n^{2}|\eta_{n}|^{2}\,=\,\sum_{n=1}^{N}n^{1+\alpha}n^{1-\alpha}|\eta_{n}|^{2}$
and then, using the Lemma in p. 101 of [12], we see that
$\sum_{n=N}^{\infty}n^{1-\alpha}|\eta_{n}|^{2}\,=\,\operatorname{O}(N^{-\alpha}),\,\,\text{as $N\to\infty$}\,\,\,\Leftrightarrow\,\,\,\sum_{n=1}^{N}n^{2}|\eta_{n}|^{2}\,=\,\operatorname{O}(N),\,\,\text{as $N\to\infty$}.$
This, (3.1), and Theorem A prove part (v) of Theorem 4. The “little oh”  analogues of the results we have just used give part (vi).
Parts (v) and (vi) of Theorem 3 follow from parts (v) and (vi) of Theorem 4 and the fact that $A^{2}_{\alpha}=\mathcal{D}^{2}_{\alpha+2}$ (see Lemma 3. 10 of [45]).
Remark 2.
For $1<\alpha\leq 0$, the sequences $(\eta)$ for which $\mathcal{R}_{(\eta)}$ is bounded (compact) on $\mathcal{D}^{2}_{\alpha}$ are characterized in Theorems 1 and 2 of [8].
4. Necessary conditions
In this section we are going to prove part (a) of theorems 1 and 2 and parts (iii) and (iv) of theorems 3 and 4.
We are going to use the following result.
Proposition 2.
For $N\,=\,2,3,4,\dots$, set $a_{N}\,=\,1-\frac{1}{N}$ and
(4.1) $\varphi_{N}(z)\,=\,\sum_{n=N}^{2N-1}\frac{1}{\sum_{k=1}^{n}ka_{N}^{k}}z^{n}\,\quad z\in\mathbb{D}.$
Then
(4.2) $\|\varphi_{N}\|_{H^{1}}\lesssim\frac{\log N}{N^{2}}.$
Proof. Notice that
where
Now, the sequence of Taylor coefficients of $\psi_{N}$ is decreasing. Then using [35, Theorem 1. 1], we obtain
$\|\varphi_{N}\|_{H^{1}}\,=\,\|\psi_{N}\|_{H^{1}}\,\asymp\,\sum_{n=0}^{N-1}\frac{1}{(n+1)\sum_{k=1}^{n+N}ka_{N}^{k}}\,\leq\,\sum_{n=0}^{N-1}\frac{1}{(n+1)\sum_{k=1}^{N}ka_{N}^{k}}\,\asymp\,\frac{\log N}{N^{2}}.$
$\Box$
For $1\leq p<\infty$ and $N\,=\,2,3,4,\dots$ set
(4.3) $f_{N,p}(z)\,=\,\frac{1}{N^{2-\frac{1}{p}}}\frac{a_{N}z}{(1-a_{N}z)^{2}}\,=\,\frac{1}{N^{2-\frac{1}{p}}}\sum_{n=1}^{\infty}na_{N}^{n}z^{n},\quad z\in\mathbb{D}.$
Proof ofTheorem 1 (a) and Theorem 2 (a). Assume that $1<p<\infty$ and $\mathcal{R}_{(\eta)}\in\mathcal{B}(H^{p})$. As observed in Section 1 this implies that $F_{(\eta)}$ is a well defined analytic function in $\mathbb{D}$. In fact, since $F_{(\eta)}=\mathcal{R}_{(\eta)}(1)$, $F_{(\eta)}\in H^{p}$.
We have that $f_{N,p}\in H^{p}$ for all $N$ and
(4.4) $\|f_{N,p}\|_{H^{p}}\lesssim 1.$